OpenStudy (anonymous):

fx=x-5 when is x is -6

OpenStudy (anonymous):

Basically, just plug -6 into the equation as x.

OpenStudy (anonymous):

fx= -6-5 fx=-11

OpenStudy (anonymous):

ok thans

OpenStudy (anonymous):

No problem :D

OpenStudy (anonymous):

i have something relted to ti can u help me

OpenStudy (anonymous):

Umm..sure

OpenStudy (anonymous):

f(x)=3x-4 domain=(-5,-3,-1,1,3)?

OpenStudy (anonymous):

how do i do this problem

OpenStudy (anonymous):

What are you solving for?

OpenStudy (anonymous):

instruction say : find the range of the function defined by the given equation write 5 ordered pairs that belong to the function

OpenStudy (anonymous):

i know this might b alont of work but is this problem where i use x and y chart?